Maple V Release 5: A Quick Reference

Prepared by: Douglas Meade Department of Mathematics University of South Carolina

April 1998, updated for Release 5 (February 1997, updated for Release 4) (January 1995, original version for Release 3)

Symbols and Abbreviations

Symbol  Description  Example

:=                 assignment  f := x^2/y^3;

;

terminate command; display result

                                                             int( x^2, x );
:

terminate command; hide result

                                                             int( x^2, x ):
..

                   specify a range or interval  plot( t*exp(-2*t), t=0..3 );
{ }

set delimiter (a set is an unordered list)

                                                             { y, x, y };
[ ]

list delimiter (lists are ordered)

                                                             [ y, x, y ];
%

refers to previous result (percent) Note: Was " in previous releases

                                                             Int( exp(x^2), x=0..1 ):
                                                             % = evalf( % );
"  "
(see ?strings)

string delimiter (double quote)

Note: New in Release 5

                                                             plot( sin(10*x) + 3*sin(x), x=0..2*Pi,
                                                                   title="An interesting plot" );

`` (see ?names)  name delimiter (back quote)  ` A name` := ` This is a name.` ;

´´ (see ?uneval) delayed evaluation (single quote)

                                                             x := ´ x´ ;
->

mapping (procedure) definition f := (x,y) -> x^2*sin(x-y);

                                                             f(Pi/2,0);
@

composition operator (cos@arcsin)(x);

@@

repeated composition operator

                                                             (D@@2)(ln);

Mathematical Operations, Functions, and Constants

Symbol  Description  Example

+, -, *, /, ^

add, subtract, multiply, divide, power

                                                             3*x^(-4) + x/Pi;
sin, cos, tan,
cot, sec, csc

                   trigonometric functions  sin( theta-Pi/5 ) - sec( theta^2 );
arcsin, arccos,
arctan, arccot,
arcsec, arccsc

inverse trigonometric functions

                                                             arctan( 2*x );

exp

exponential function

                                                             exp( 2*x );
ln

natural logarithm

                                                             ln( x*y/2 );
log10

common logarithm (base 10)

                                                             log10( 1000 );
abs

absolute value

                                                             abs( (-3)^5 );

sqrt               square root  sqrt( 24 );

!                  factorial  k!;

=, <>, <, <=, >, >=

equations and inequalities

Note: E no longer exists; use exp(1)

                                                             diff( y(x), x ) + x*y(x) = F(x);
                                                             exp(Pi) > Pi^exp(1);
Pi, I

π, i (mathematical constants) Note: Maple is case-sensitive

                                                             exp( Pi*I );
infinity

infinity

                                                             int( x^(-2), x=1..infinity );

\underline{NOTES}:

• The document is also available on the World Wide Web; the Universal Resource Locator is

    http://www.math.sc.edu/~meade/maple/maple-ref/

• Please send comments, corrections, and suggestions for improvements to meade@math.sc.edu.